14. Bromine iodide decomposes according to the equation: 2Brl(g) ? Br2(g)+12(g) Keq-00156 A 0.660 mol sample of Brl was injected into a 2.00 L reaction vessel. and 12 at equilibrium. Calculate the concentration of Brl, Br2,
Answer #1

2 BrI (g) \rightleftharpoons Br2 (g) + I2 (g)

Initial

[BrI] = 0.660 / 2.00 = 0.330 M

[Br2] = [I2] = 0.000 M

At equilibirum,

[BrI] = ( 0.330 - 2x ) M

[Br2] = [I2] = x M

Equilibrium constant expression can be written as,

Keq = [Br2][I2] / [BrI]2

0.0156 = x * x / ( 0.330 - 2x)2

0.0156 * ( 0.1089 + 4x2 - 1.32 x ) = x2

0.00170 + 0.0624 x2 - 0.0206 x = x2

0.9376 x2 + 0.0206 x - 0.00170 = 0

Applying quadartic equation to solve x,

x = [ - 0.0206 +/- sqrt((0.0206)2 - 4 * 0.9376 * ( - 0.00170 )) ] / ( 2 * 0.9376 )

x = 0.0330 M

Therefore,

at equilibrium,

[BrI] = 0.330 - 2 * 0.0330 = 0.264 M

[Br2] = [I2] = 0.0330